Show that x = y = z+1 is a secant line of the sphere x^2 +y^2 +z^2 −x−y+z−1 = 0.
Also find the intercept made by the sphere on the line.
1
Expert's answer
2020-05-13T20:10:14-0400

The line is secant of the sphere if the line has two intersection points with the sphere. Let us find these points. Let us substitute "y = x = z+1" in the equation of sphere:

"x^2+x^2+(x-1)^2 - x - x + (x-1 )-1 = 0." Therefore, "3x^2-3x-1=0."

Next, we calculate the discriminant of the equation:

"D^2 = b^2-4ac =" 21. We see that "D^2>0," so the equation has two roots and the line has two points of intersection with the sphere, because every root corresponds to one "y" and "z" value.

Now let us determine the roots:

"x_1 = \dfrac{3+\sqrt{21}}{6}, \;\; x_2 = \dfrac{3-\sqrt{21}}{6}."

Next we calculate "y" and "z" coordinates:

"y_1=x_1 = \dfrac{3+\sqrt{21}}{6}, \;\; y_2= x_2 = \dfrac{3-\sqrt{21}}{6}; \\
z_1 = x_1-1 = \dfrac{-3+\sqrt{21}}{6}, \;\; z_2=x_2-1 = \dfrac{-3-\sqrt{21}}{6}."

Therefore, points are

"\left( \dfrac{3+\sqrt{21}}{6}, \dfrac{3+\sqrt{21}}{6}, \dfrac{-3+\sqrt{21}}{6} \right) \mathrm{and} \left( \dfrac{3-\sqrt{21}}{6}, \dfrac{3-\sqrt{21}}{6}, \dfrac{-3-\sqrt{21}}{6} \right)."

We may also calculate the length of a segment between these points:

"l = \sqrt{\left(x_2-x_1 \right)^2 +\left(y_2-y_1 \right)^2 + \left(z_2-z_1 \right)^2} = \sqrt{7}."